OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle: cos25°cos15°-sin25°sin15°

OpenStudy (anonymous):

@Jhannybean

OpenStudy (shamim):

the formula is\[\cos(A+B)=cosAcosB-sinAsinB\]

OpenStudy (anonymous):

how do i know to use the cos formula? and not the sine formula?

OpenStudy (shamim):

because ur given equation is\[\cos25\cos15-\sin25\sin15=\cos(25+15)=\cos40\]

OpenStudy (shamim):

medal

OpenStudy (anonymous):

??

OpenStudy (shamim):

if u r satiisfied then give me a medal

OpenStudy (shamim):

just click best response

OpenStudy (anonymous):

thanks for the help